OpenStudy (anonymous):

Use spherical coordinates to find a parametrization of the quarter-sphere x2+y2+z2=a2, z ≥ 0, x ≥ 0

OpenStudy (anonymous):

rho,theta,phi

OpenStudy (anonymous):

\[\huge\rho = x^2+y^2+z^2 \] \[\huge\theta= arctan\left(\frac{y}{x}\right) \] \[\huge \phi = arccos\left(\frac{z}{\sqrt{x^2+y^2+z^2}} \right)\]